how do i find the time it takes a ball dropped from 10ft. to hit the ground?
Answer by scott8148(6628)

gravitational acceleration is 32ft/sec^2

since the ball starts from rest and free falls ___ 10 = (32/2) t^2

5/8 = t^2

t = sqrt(5/8)